Determine the weight of the cube if the weight of the paraffin oil in the cube is 79.2 pound/ft3 and edge length of the cube
is 6 in.

Answer by mananth(16946)

Determine the weight of the cube if the weight of the paraffin oil in the cube is 79.2 pound/ft3 and edge length of the cube
is 6 in.
volume of cube = L^3
length = 1.2 foot
volume = (1/2 )3
= 1/8 cubic foot
1 cu,ft = 79.
1/8 ft^3= 1/8 * 79.2
=9.9 pounds